JET-COOLED LASER SPECTROSCOPY OF CYCLOHEXOXY RADICAL

LILY ZU, \underline{JINJUN LIU}, GYÖRGY TARCZAY,^{a} PATRICK DUPRÉ and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University,120 W. 18th Avenue, Columbus, Ohio 43210.

The laser induced fluorescence(LIF), and laser excited, dispersed fluorescence(DF) spectra of the cyclohexoxy free radical have been observed under free-jet cooling conditions for the B̃ − X̃ electronic transition. In the LIF excitation spectrum, two different types of expansions, characterized by rotational temperatures of 1 and 100 K, have demonstrated that a number of lines, termed hot bands, either nearly or entirely disappear at the lowest jet temperature. The cold spectral bands have two distinct types of structure in rotationally resolved high-resolution spectra. Correspondingly, two distinct sorts of structure are observed in the DF spectra. LIF transitions to both a′ and a′′ vibrational levels in the B̃ state are observed and allowed due to a substantial pseudo-Jahn-Teller effect in the X̃ state. We assign hot bands to transitions to the origin and excited vibrational levels of the B̃^{2}A′ state from the vibrationless level of the low-lying Ã electronic state. Analyses of the spectra yield vibrational frequencies for the X̃, Ã, and B̃ states as well as the energy separations of their vibrationless levels. Although 5 conformers of cyclohexoxy are possible, it appears that all presently observed spectral bands can be accounted for by a single conformer. A more complete rotational analysis is now in progress.

aPresent address: Department of General and Inorganic Chemistry, Eötvös University, P. O. Box 32, H-1518 Budapest 112, Hungary